### v3

# Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image]

Abstract
Urban causal research is essential for understanding the complex, dynamic processes that shape cities and for informing evidence-based policies. However, current practices are often constrained by inefficient and biased hypothesis formulation, challenges in integrating multimodal data, and fragile experimental methodologies. Imagine a system that automatically estimates the causal impact of congestion pricing on commute times by income group or measures how new green spaces affect asthma rates across neighborhoods using satellite imagery and health reports, and then generates comprehensive, policy-ready outputs, including causal estimates, subgroup analyses, and actionable recommendations. In this Perspective, we propose UrbanCIA, an LLM-driven conceptual framework composed of four distinct modular agents responsible for hypothesis generation, data engineering, experiment design and execution, and results interpretation with policy insights. We begin by examining the current landscape of urban causal research through a structured taxonomy of research topics, data sources, and methodological approaches, revealing systemic limitations across the workflow. Next, we introduce the design principles and technological roadmap for the four modules in the proposed framework. We also propose evaluation criteria to assess the rigor and transparency of these AI-augmented processes. Finally, we reflect on the broader implications for human–AI collaboration, equity, and accountability. We call for a new research agenda that embraces LLM-driven tools as catalysts for more scalable, reproducible, and inclusive urban research. Equal contribution.
[Uncaptioned image] Reimagining Urban Science: Scaling Causal Inference with Large Language Models
Yutong Xia[Uncaptioned image][Uncaptioned image]  Ao Qu[Uncaptioned image]  Yunhan Zheng[Uncaptioned image]  Yihong Tang[Uncaptioned image]  Dingyi Zhuang[Uncaptioned image] Yuxuan Liang[Uncaptioned image]  Shenhao Wang[Uncaptioned image]  Cathy Wu[Uncaptioned image]  Lijun Sun[Uncaptioned image] Roger Zimmermann[Uncaptioned image]  Jinhua Zhao[Uncaptioned image] [Uncaptioned image]MIT  [Uncaptioned image]NUS  [Uncaptioned image]McGill  [Uncaptioned image]HKUST (GZ)  [Uncaptioned image]UFL {yutongx,qua,yunhan,dingyi,cathywu,jinhua}@mit.eduyihong.tang@mail.mcgill.ca lijun.sun@mcgill.cayuxliang@outlook.comshenhaowang@ufl.edudcsrz@nus.edu.sg
Cities have the capability of providing something for everybody, only because, and only when, they are created by everybody.
Jane Jacobs
1 Introduction
Cities are complex systems shaped by interconnected forces, from transportation congestion and housing affordability to climate resilience and equitable public services Ritchie et al. (2024); United Nations (2015); Xia et al. (2022). Scientific progress in urban research equips policymakers, planners, and communities with the knowledge to design evidence-based interventions that shape urban futures Batty (1997). At the heart of this scientific effort is the pursuit of causal understanding: uncovering how policies, infrastructures, and behaviors interact to produce measurable impacts Baum-Snow and Ferreira (2015). Urban causal research typically follows a structured pipeline: beginning with a hypothesis, integrating diverse data sources, selecting appropriate methods, and interpreting results to inform decisions. For example, in assessing New York City’s congestion pricing policy News (2024), scientists may hypothesize that its impact on commute times differs by neighborhood income levels. This triggers a series of steps: collecting multimodal data such as subway logs, taxi trips, social media, and census records; applying methods like difference-in-differences with fixed effects Hernan and Robins (2020). The final analysis reveals not only whether commute times decreased, but whether the policy disproportionately benefits certain socioeconomic groups. Such approaches have also enabled studies on the effects of high-speed rail Heuermann and Schmieder (2019), electric vehicle infrastructure Zheng et al. (2024a), and vaccine mandates Karaivanov et al. (2022).
However, this process remains labor-intensive, fragile, and unevenly accessible. We identify three major challenges: 1) Hypothesis discovery is slow and biased: research questions are often manually crafted and shaped by disciplinary precedent or funding priorities, overlooking emerging issues and underrepresenting smaller cities or informal settlements Bell and Jayne (2009); Müller et al. (2025); 2) Data complexity impedes integration: although urban environments generate rich and diverse data streams (e.g., satellite imagery, mobility logs, text reports), integrating them into causal studies remains technically demanding, leading many studies to rely primarily on structured tabular datasets Raghavan et al. (2020); Hong et al. (2023); 3) Experimentation is expensive and risk-prone: causal inference depends on stringent assumptions, such as valid counterfactuals or exogenous variation, and many studies are abandoned when these conditions fail, discouraging exploration of high-risk, high-reward policy questions Baum-Snow and Ferreira (2015); Matthay and Glymour (2022). Taken together, these challenges underscore the need for more systematic, data-aware, and technology-enabled approaches to causal inference in urban contexts. The structural limitations of current urban causal research raise a timely question: Can artificial intelligence help expand the scope, efficiency, and inclusivity of urban scientific discovery?
Recent advances in Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) and Multimodal Large Language Models (MLLMs) Wu et al. (2023) offer new possibilities for reimagining how urban causal research is conducted. LLMs exhibit strong capabilities in language understanding Radford et al. (2018), reasoning Wei et al. (2022); Jaech et al. (2024), planning Huang et al. (2022), code execution Chen et al. (2021), and tool use Schick et al. (2023); Qin et al. (2024). MLLMs extend these functions across diverse data modalities such as imagery Liu et al. (2023a), audio Chu et al. (2023), graph Tang et al. (2024), structured charts Han et al. (2023), and unstructured reports Li et al. (2024a), lowering the barriers to engaging with complex urban data at scale Liang et al. (2025). Despite these technical advances, their application to urban causal research remains fragmented, often limited to isolated tasks such as data cleaning, code generation, or report drafting. Recent developments in agent-based LLM systems Lu et al. (2024); Schmidgall et al. (2025); Gottweis et al. (2025) suggest a promising direction: not using LLMs as end-to-end black boxes, but as specialized components within a coordinated reasoning system.
Figure 1: Overview of the paper’s structure and logic. A layered structure from review to framework and broader discussion. Arrows indicate how insights from one stage inform the next.
In this Perspective, we argue that LLMs offer a path to reimagine how we ask questions, extract evidence, and generate causal knowledge in cities - at scale, and with broader reach. To explore this potential, we introduce UrbanCIA, a flexible and modular framework for Urban Causal Inference with Agents. UrbanCIA is a conceptual framework that supports the full urban causal pipeline: surfacing overlooked questions, assembling multimodal data, evaluating methodological assumptions, and generating interpretable outputs. Compared to conventional manual workflows and isolated AI-assisted tools, UrbanCIA offers a structured, extensible, and interpretable framework that supports the full urban causal inference process while preserving space for human judgment, domain expertise, and ethical deliberation.
The outline of this paper are shown in Figure 1. Specifically, Section 2 presents a systematic review of urban causal research, identifying key gaps in data, methods, and geographic coverage. Section 3 introduces UrbanCIA, a modular agent-based framework that supports all stages of causal inference, with a concrete multi-agent system use case. We also discuss potential challenges and propose evaluation methods. Section 4 reflects on how UrbanCIA enables broader participation from civic groups, governments, and urban communities, and offer a perspective on the human-AI collaboration paradigm. Finally, Section 5 summarizes the paper and outlines future directions. Our contributions are summarized as follows:
  • A Comprehensive Systematic Review. We conduct a large-scale empirical analysis of causal inference studies in urban research, covering over 2,400 articles across a decade from a flagship journal (i.e., Cities), identifying key trends and gaps in the use of causal inference under the urban context.
  • A Modular Framework. We introduce UrbanCIA, a flexible, AI-augmented pipeline that supports all stages of causal analysis through modular agents spanning hypothesis generation, data preparation, experimental execution, and report writing, demonstrated through a concrete multi-agent system use case.
  • A Multi-Dimensional Evaluation Protocol. We examine methodological, ethical, and deployment challenges, and propose metrics for assessing the rigor, novelty, and generalizability of AI-generated causal research.
  • Towards Inclusive Urban Science. We show how UrbanCIA reduces the entry barriers to rigorous urban causal research, allowing broader participation by civic organizations, municipal governments, and community groups - advancing the vision of a city “created by everybody” Jacobs (1961).
2 Landscape of Causal Inference in Urban Studies
This section provides a quantitative overview of the field based on a comprehensive empirical analysis. It identifies key trends in causal inference research across urban topics, data usage, methodological approaches, and geographic focus. These findings lay the foundation for designing UrbanCIA, our proposed AI framework for automated causal research, ensuring it is aligned with actual research needs, practices, and gaps (see Appendix C for methodological details).
2.1 Why a Systematic Review?
- Mapping the current landscape to inform future design. To develop an effective AI system for causal inference in urban research, it is critical to first understand how such research is currently conducted. Urban systems are inherently complex, characterized by dynamic, interlinked processes spanning infrastructure, policy, society, and environment. Identifying causal relationships within such systems is essential for evaluating the impact of interventions, but the methodological choices, data modalities, and research foci within this domain remain unevenly distributed and often under-documented.
Figure 2: Rising adoption of causal inference in urban research. Trends in publication volume (blue bars) and causal inference usage (red bars) in Cities from 2012 to 2021. The dotted line marks the growing share of studies employing causal methods over time.
2.2 How We Reviewed It?
- Data Collection. We conducted a systematic review of 2,428 peer-reviewed articles published in Cities, a leading journal in urban research, spanning a 10-year period from 2012 to 2021. The related publication trend of urban causal inference studies is shown in Figure 2. Using a pipeline combining state-of-the-art LLMs, i.e., GPT-4o-mini, GPT-4o, and GPT-4.5 Jaech et al. (2024), and classical machine learning techniques, followed by expert human validation, we identified 219 articles that employed formal causal inference methodologies in their urban studies.
- Analysis Procedure. For each paper, we extracted metadata along five dimensions: Urban Topic (e.g., housing, transportation), Geographic Focus (i.e., continent and country), Specific Data Used (e.g., panel data, satellite imagery), Causal Methodology (i.e., method family and technique), and Reproducibility (i.e., availability of code and data). The results from the analysis are presented through visualizations in Figure 3 (i.e., distributions of the investigated dimensions).
Figure 3: Key gaps in urban causal inference research. (a) Distribution of 219 studies across topics, regions, data types, experimental designs, and methods, revealing narrow topical coverage and methodological homogeneity. (b) Global disparities, with over 75% of studies from China, the United States, and Spain. (c) Limited use of multimodal data, with 64% of studies relying solely on structured tabular sources.
2.3 What We Found?
- Key trends and empirical patterns. Our systematic review reveals a field that is gaining momentum but remains uneven and constrained in several dimensions. While it has seen growing adoption over the past decade, it is still marked by geographical imbalances, a heavy reliance on structured data, limited methodological diversity, and poor reproducibility, detailed as follows.
  • A Growing Emphasis on Causal Research. Causal inference has attracted increasing attention from urban science scholars over the past decade. As shown in Figure 2, in 2012, only 2 out of 67 studies (3.0%) in Cities used causal inference methods. By 2020, this proportion rose to 64 out of 421 studies (15.2%). From the growing trend shown in the figure, we can see a clear shift toward evidence-based evaluation in urban policy research.
  • Geographical Imbalances. As indicated in Figure 3b, urban causal research is heavily concentrated in a few countries. The top three, China (112 studies), United States (40), and Spain (16), together account for over 75% of the total causal inference corpus. In contrast, entire regions such as Sub-Saharan Africa, South Asia, and Latin America remain significantly underrepresented. Even countries undergoing major urban transitions, such as India, Brazil, and South Africa, each have fewer than 10 studies. These imbalances raise concerns about the generalizability of findings and underscore the need for scalable solutions that lower entry barriers for underrepresented regions.
  • Overreliance on Structured Data. The majority of studies rely on structured tabular datasets, including census records, economic indicators, and administrative databases. Based on our classification in Figure 3c, 64% of studies use structured data and time series panel data exclusively. Only 12% incorporate geospatial or vector data, 4% utilize image or remote sensing data, and 14% analyze textual or document-based data. This underutilization of multimodal data reflects analytical limitations, an area where AI systems can offer significant advances.
  • Limited Methodological Diversity. As shown in Figure 3a, methodological approaches are dominated by conventional econometric techniques. Among the 219 papers, 38.4% use Fixed Effects Panel Regression, 23.7% apply Difference-in-Differences (DiD), 17.4% use Instrumental Variables (IV), and 11.9% rely on Propensity Score Matching (PSM). In contrast, machine learning-based methods such as causal forests appear in fewer than 3% of studies. Bayesian models and structural equation modeling are similarly rare, despite their strengths in modeling latent and complex constructs.
  • Lack of Transparency and Reproducibility. Among the 219 causal inference papers reviewed, none provided open access to their codebase, and fewer than 15% offered accessible data. This lack of transparency significantly hinders reproducibility and restricts opportunities for secondary analysis or AI-assisted meta-learning.
2.4 Insights from Empirical Landscape
The findings from our systematic review point to several critical directions for designing future AI-augmented systems for urban causal research. First, broader topical coverage is needed. Current studies are concentrated in a narrow set of urban domains, leaving governance, inequality, and service delivery underexplored. Second, systems should natively support diverse and unstructured data formats, such as geospatial vectors, imagery, and textual records, which are currently underutilized but essential for capturing complex urban phenomena. Third, the heavy reliance on conventional econometric models suggests a need for more flexible methodological toolkits that can adapt to data properties and research questions. Fourth, the lack of transparency and open access underscores the importance of embedding reproducibility and accountability into system design (e.g., generating logs, versioned outputs, and shareable code). Finally, geographic imbalances highlight the need for systems that lower technical entry barriers, enabling participation from data-scarce or underrepresented regions. These implications provide foundational guidance for future tools that seek to make urban causal inference more inclusive, transparent, and methodologically robust.
3 UrbanCIA framework
This section proposes UrbanCIA, a modular and flexible framework for urban causal inference. Specifically, we begin by identifying key barreiers of current manual workflows (Figure 4, left) and then highlight the capabilities of LLMs/MLLMs that make automation feasible (Figure 4, right). Building on these two foundations, we introduce a modular, agent-based framework that structures the causal inference process into four components (Figure 4, center), each supported by different operational modes and LLM capabilities. We then present a concrete multi-agent use case to illustrate the system in action. Finally, we discuss key challenges and outline an evaluation strategy for UrbanCIA.
3.1 Motivation
Why automation is needed?
Urban causal inference requires coordination across multiple stages, from defining questions to delivering policy-relevant insights. While our systematic review in Section 2 reveals increasing interest in causal methods, it also uncovers persistent limitations in current practice, i.e., overreliance on structured data, limited methodological diversity, and a lack of reproducibility. These issues stem in part from how each stage of the causal workflow is currently executed (Figure 4, left). In Phase 1, generating hypotheses often relies on subjective intuition or delayed academic cycles, making it difficult to identify timely and novel research questions. In Phase 2, assembling urban data is frequently repetitive and labor-intensive, requiring manual retrieval, cleaning, and integration of multi-scource (e.g., traffic sensors, social media, and environmental data) multimodal (e.g., image, spatial time series, and text), multiresolution (e.g., different smallest units or time intervals) datasets Zou et al. (2025); Zhang et al. (2024b). These processes are time-consuming and difficult to generalize across cities. Phase 3 focuses on executing causal inference experiments. Here, researchers must not only select the most appropriate estimation strategy (e.g., PSM, DiD, or IV) but also ensure alignment with the assumptions and structure of the available data, which often demands substantial methodological expertise. Finally, in Phase 4, drawing policy insights involves synthesizing findings into accessible narratives for decision-makers, often requiring interdisciplinary translation that is rarely systematized. These limitations motivate the need for a unified, modular framework that supports automation across all four phases while maintaining scientific rigor, flexibility, and interpretability.
Figure 4: From limitations to design: motivating the UrbanCIA framework.. The challenges of manual urban causal workflows (left), the modular design of UrbanCIA offering agents with multiple operational modes for each stage (middle), and the capabilities of LLMs/MLLMs that can support distinct stages, denoted by colored checkmarks (right).
What LLMs make possible?
Recent advances in large language and multimodal models (LLMs and MLLMs) have introduced new possibilities for automating causal inference workflows in urban research. We organize their relevant capabilities into two core categories: knowledge reasoning and procedural execution, which correspond closely to the intellectual and operational demands of the UrbanCIA framework.
Knowledge reasoning encompasses the ability to interpret contextual information, generate novel ideas, and communicate structured findings. This includes semantic understanding of spatial, temporal, and causal concepts Zhao et al. (2024); Zhang et al. (2024b), creative hypothesis formation from literature, data, or policy narratives Si et al. (2024); Chakrabarty et al. (2024), and the capacity to articulate experimental insights into interpretable outputs such as research summaries or policy recommendations Chia et al. (2024); Baek et al. (2024).
Procedural execution refers to the model’s capacity to interact with external tools and data ecosystems, generate executable code, and perform end-to-end causal workflows. LLMs and MLLMs can retrieve urban datasets via APIs Yang et al. (2024b), align multimodal inputs Gupta and Kembhavi (2023), and implement causal estimation methods by synthesizing and running code across statistical and machine learning libraries Qin et al. (2024); Li et al. (2023); Huang et al. (2024b).
These capabilities provide the foundation for the agents within UrbanCIA, enabling both cognitive support (e.g., hypothesis refinement, narrative construction) and operational automation (e.g., data acquisition, model execution) across the entire urban causal inference pipeline, as illustrated in the right portion of Figure 4.
3.2 Framework structure and agent modes
Motivated by the above-mentioned limitations and emerging LLM/MLLMs’ capabilities, we propose UrbanCIA —a modular framework that structures urban causal inference into functional phases, with LLM agents assigned scoped roles and distinct modes of interaction. As shown in Figure 4 (middle), UrbanCIA is a flexible framework composed of four functional phases. Note that each phase can be supported by one or more agents, and agents can also operate across multiple phases as needed. Instead of a fixed pipeline, the framework provides a modular structure that adapts to varying data availability, task complexity, and automation needs. This flexibility allows UrbanCIA to support diverse use cases, including real-time policy decisions and data-scarce environments. The agent setup described here is one example; a concrete multi-agent instantiation is introduced in Section 3.3. We then describe each phase with different mode in detail below.
Hypothesis Generation Agents.
The first phase of the causal workflow involves identifying meaningful and timely causal questions (i.e., hypothesis) rooted in urban phenomena. Traditionally, this step relies heavily on domain expertise, often guided by literature reviews or retrospective interpretations of policy outcomes—an approach that can introduce cognitive bias and limit scalability. To address this, UrbanCIA supports multiple operational modes, each tailored to different types of input sources for hypothesis generation.
  • In the theory-driven mode, agents assist in refining or critiquing human-provided hypotheses by referencing academic literature and established theoretical frameworks.
  • The data-driven mode identifies anomalies or shifts in sensor-based observations (e.g., spikes in traffic congestion or air pollution) that may point to candidate interventions or underlying causal mechanisms.
  • The practice-driven mode draws inspiration for causal questions from real-world information sources, including policy documents, urban news feeds, and social media narratives, grounding hypotheses in ongoing societal dynamics.
These modes can be invoked independently or in combination, depending on the research objective and the nature of available inputs.
Urban Data Agents.
The second phase focuses on constructing datasets for downstream causal experiments, often from multi-source, multi-modal, and multi-resolution urban data. To accommodate varying levels of data accessibility and completeness, agents in this phase may operate under three distinct modes.
  • In the user-uploaded data mode, agents process datasets provided directly by researchers, handling resolution alignment and basic preprocessing. This mode is suitable when users bring domain-specific or proprietary data into the workflow.
  • In the web-based retrieval mode, agents autonomously identify and collect relevant datasets from open data platforms, online repositories, or real-time APIs, guided by the variables required to test a given hypothesis. Examples of such sources include Google Street View (GSV) [^1] https://www.google.com/streetview, OpenStreetMap (OSM) [^2] https://www.openstreetmap.org, and official government portals such as NYC Open Data [^3] https://opendata.cityofnewyork.us/ and Singapore Open Data [^4] https://data.gov.sg/.
  • In the simulation-based data mode, agents generate synthetic data to fill the missing data to support counterfactual reasoning, robustness testing, and early-stage experimentation when real-world data are costly, sparse, or unavailable. Recent works leverage LLMs to simulate structured urban behaviors—such as generating vehicle trajectories from textual descriptions Yang et al. (2025) or modeling human mobility using personas and routine profiles Ju et al. (2025); Wang et al. (2024). These synthetic data environments can approximate real-world urban conditions, offering a flexible proxy for evaluating causal questions when actual interventions are infeasible.
CI Experiment Agents.
Once the dataset has been prepared, agents in this phase carry out causal inference experiments to estimate the effects of candidate interventions. Depending on the structure of the data, the nature of the hypothesis, and the complexity of the treatment assignment, different estimation modes may be activated. In many cases, agents implement these experiments by generating executable code, typically in Python or R, tailored to the selected method, and then running the code to produce effect estimates. Commonly used libraries include DoWhy for end-to-end causal inference workflows Sharma and Kiciman (2020), EconML for machine learning-based treatment effect estimation Battocchi et al. (2019), and CausalML for uplift modeling and heterogeneous effect analysis Chen et al. (2020).
  • In the randomized designs mode, agents apply experimental frameworks such as RCTs Hernan and Robins (2020) or their approximations when random assignment is feasible, observed, or can be simulated.. In urban research, true randomization is rarely possible due to ethical, logistical, and political constraints. Thus, this mode is most often used in conjunction with simulation-based data generated in the previous phase. When treatment and control groups can be synthetically constructed, agents can estimate causal effects under strong identification assumptions.
  • The quasi-experimental methods mode is suitable when natural or institutional variation can be leveraged as a source of exogenous treatment. Agents operating in this mode identify applicable designs such as DiD, PSM, IV, or SCM Hernan and Robins (2020), depending on the data structure and policy context.
  • In the ML-based estimation mode, agents use data-driven models to estimate treatment effects in high-dimensional or heterogeneous settings. These may include causal forests Ronco et al. (2023); Ito et al. (2024), deep neural networks for counterfactual inference Shalit et al. (2017); Louizos et al. (2017); Shi et al. (2019), and their recent extensions to graph-structured data Wein et al. (2021); Xia et al. (2023). This mode is well suited for settings with treatment effect heterogeneity or limited structural assumptions.
By supporting diverse estimation strategies and automating their execution, this phase enables agents to conduct causal inference across varied urban scenarios, from controlled interventions to observational settings.
Evaluation Agents.
Following causal estimation, agents in this final phase perform critical evaluations of the experimental results and help translate them into actionable insights. Depending on the research context, this process can involve technical validation, interpretability enhancement, and policy communication.
  • In the Robustness Test mode, agents assess the sensitivity of causal estimates to modeling choices, alternative specifications, and sample heterogeneity. This includes placebo tests Eggers et al. (2024), falsification analyses, and subset re-estimation to verify whether observed effects persist under reasonable perturbations.
  • The Validity Check mode focuses on evaluating identification assumptions and threats to internal or external validity. Agents may examine covariate balance, assess pre-treatment trends (e.g., in DiD settings), or conduct refutation tests to identify potential confounders and post-treatment bias Hernan and Robins (2020).
  • In the Policy Narrative mode, agents assist in constructing interpretable summaries and decision-relevant narratives grounded in experimental findings. This includes generating textual explanations, translating results into human-readable formats, and mapping causal insights to policy goals and stakeholder concerns.
Together, these modes ensure that causal insights are not only statistically sound but also communicable and useful for real-world urban decision-making. Notably, we restate the flexibility of agent coordination across the framework. Each phase can be supported by one or more agents, and agents may operate across multiple phases as needed. The workflow is not strictly linear: for instance, an experiment agent may determine that an IV is required, prompting the data agent to identify a suitable instrument; or, if a PSM method is selected, the evaluation agent may test for parallel trends to confirm its validity. This modular and adaptive architecture enables agents to coordinate iteratively, refining the pipeline based on evolving task demands. We illustrate this flexibility in the following use case (Section 3.3).
3.3 Multi-agent system example
This section presents a concrete instantiation of UrbanCIA through a multi-agent system (MAS) example, illustrating how agents collaborate across the causal workflow in a realistic research scenario. We focus on a case study investigating the impact of congestion pricing on low-income commuters in New York City (NYC), as shown in Figure 5. This example demonstrates one possible path through the framework: Practice-Driven + Theory-Driven $\rightarrow$ Web-Based Retrieval $\rightarrow$ Quasi-Experimental Methods $\rightarrow$ Validity Check + Policy Narrative + Robustness Test. The system includes seven specialized agents—Urban Scientist, Reader, Data Engineer, Validator, Data Scientist, Experimenter, and Writer—each responsible for a distinct function in the pipeline. Detailed roles and responsibilities are described in Appendix D.
Figure 5: An example of the UrbanCIA system in practice. Phase 1 Hypothesis Generation: The Reader gathers policy news, social media posts, and academic papers, proposes a preliminary hypothesis, and refines it with the Urban Scientist while identifying the required data types. Phase 2 Dataset Preparation: The Data Engineer collects and preprocesses multimodal urban data. The Validator checks the validity of the resulting dataset. Phase 3 Experimental Design & Execution: The Experimenter runs exploratory regressions and conducts experiments under the design proposed by the Data Scientist. The Validator checks the robustness of the results. Phase 4 Results Interpretation: The Writer synthesizes results into a structured report in collaboration with the Data Scientist and Urban Scientist. The Validator checks the narrative alignment.
Phase 1: Hypothesis Generation.
This phase includes a continuous, two-stage process to identify emerging urban challenges and translate them into researchable questions. In this example, the system monitors public discourse surrounding NYC’s recently announced congestion pricing policy. A Reader agent continuously collects multimodal information from social media platforms (e.g., X), online news sources, and academic search engines. Examples include news headlines such as ‘New York to become first US city to have congestion charge” News (2024), and social media posts expressing concerns like “\$9 just to drive into Manhattan? What about working-class folks?” These inputs are combined with academic literature retrieved from sources such as Google Scholar on prior congestion pricing implementations in cities like Stockholm Chronopoulos (2012); Börjesson et al. (2012). It proposes a preliminary hypothesis, which is then refined by the Urban Scientist through contextual knowledge of NYC’s transit systems and demographic structure. The final causal hypothesis states: “In NYC, congestion pricing reduces commute times more for higher-income commuters than for lower-income commuters, as the latter are more likely to be priced out of driving and forced to rely on public transit options.” Relevant data sources are also identified, including subway and bus logs, taxi trips, street view images, social media posts, and census-based income records.
Phase 2: Dataset Preparation.
Once the hypothesis is defined and relevant data requirements are identified, the Data Engineer constructs a pipeline to collect, process, and integrate multimodal urban data into a structured, analysis-ready dataset. This involves cross-domain web-based retrieval of spatio-temporal, visual, and textual data Zou et al. (2025), spanning street view imagery, transit logs (subway, bus, taxi), social media posts, and census income records Zhang et al. (2024b). To extract features from heterogeneous modalities, Vision-Language Models (VLMs) Bordes et al. (2024) and the Segment Anything Model (SAM) Kirillov et al. (2023) are applied to visual inputs, while models like LLMs assist in extracting semantic attributes from text. Geo-location information is obtained either from structured metadata (e.g., geo-tags) or through named entity recognition and geocoding techniques applied to textual content Yan et al. (2024). Standard preprocessing steps—data cleaning, normalization, categorical encoding, and spatio-temporal alignment—ensure consistency across scales and formats.
The resulting dataset is evaluated by a Validator, which checks whether it satisfies core causal requirements such as temporal coverage, treatment/control group availability, and covariate completeness Bailey et al. (2024). If conditions met, it proceeds to the next phase; otherwise the system initiates an iterative refinement process, updating the data pipeline or revisiting the hypothesis as needed.
Phase 3: Experimental Design & Execution.
With a validated dataset, the Experimenter first conducts a preliminary regression analysis (e.g., OLS) to provide the Data Scientist with diagnostic insights on variable relevance and potential confounders. Guided by these results, the Data Scientist selects appropriate causal inference methods based on the data structure and hypothesis. For the NYC congestion pricing case, it adopts a DiD approach with fixed effects as the primary strategy, complemented by a SCM as a sensitivity check Hernan and Robins (2020). Key confounders, e.g., weather, route closures, and service delays, are specified, and assumptions like parallel pre-treatment trends will need to be tested. The Experimenter then implements the experimental design through code execution, including DiD and SCM estimation, subgroup analysis by income quantiles. The resulting outputs are reviewed by the Validator, who assesses the robustness (e.g., placebo tests Eggers et al. (2024)). If the results satisfy statistical and causal validity criteria, the system proceeds generation of visualizations and summary tables to the next phase; otherwise, it reverts to earlier stages for refinement.
Phase 4: Results Interpretation.
The Writer synthesizes outputs from the causal analysis in collaboration with the Urban Scientist and Data Scientist. The Data Scientist outlines methodological choices and limitations, while the Urban Scientist interprets effect sizes in the context of urban policy and equity. These perspectives are integrated into a structured report, combining visualizations, statistical justifications, and actionable insights to ensure policy relevance. The Validator then reviews the report’s narrative coherence and policy alignment. If approved, the system proceeds to the finalized reports; otherwise, earlier stages are revisited for refinement.
3.4 Challenges
We outline key challenges in realizing UrbanCIA as a scalable, policy-relevant system for causal analysis in urban environments.
Data Integration & Quality.
The first challenge arises from the inherent complexity of urban data, which are multi-source, multi-granular, and multi-modal. Specifically, urban data originate from highly heterogeneous sources, including social media, satellite imagery, mobility logs, environmental sensors, and policy documents. These datasets differ not only in modality (e.g., text, images, time series, and tabular formats) but also in granularity, ranging from neighborhood-level attributes to real-time street-level observations Zhang et al. (2024b); Zou et al. (2025). Integrating such data presents substantial challenges in preprocessing, normalization, and semantic alignment Martínez-García and Hernández-Lemus (2022); Liang et al. (2025). Recent work has explored construct urban knowledge graphs from unstructured data Liu et al. (2023b); Ning and Liu (2024), which aim to represent complex urban entities and their relationships in a structured, machine-readable form. While promising, such efforts differ from our objective: rather than building relational graphs, our framework seeks to convert multi-modal observations into structured, temporally aligned datasets that support causal inference experiments.
Maintaining Consistency in Multi-Agent Interactions.
MAS-based workflows often require agents to engage in multi-turn reasoning, where context from earlier stages (e.g., hypothesis generation) must persist throughout the pipeline. However, LLM-based agents frequently exhibit context drift, where key information from earlier prompts is forgotten or ignored in later stages Guo et al. (2024). This issue mirrors recent findings in multi-round simulations, where attention mechanisms fail to prioritize initial instructions as the context window expands Yang et al. (2024a). Proposed solutions Touvron et al. (2023) offer promising directions for improving long-range coherence. Still, systematically benchmarking and improving agent-level consistency in causal inference workflows remains an open challenge.
Aligning Agents with Domain-Specific Causal Frameworks.
While UrbanCIA leverages LLM agents for flexible reasoning, these models are not inherently grounded in established causal methodologies. Without explicit constraints, agents may produce plausible-sounding but methodologically invalid conclusions, particularly in quasi-experimental designs that require strict adherence to identification assumptions. Moreover, imposing rigid workflow templates may limit adaptability to diverse research designs and non-standard causal frameworks Schmidgall et al. (2025). Effective alignment may involve expert-annotated prompting pipelines Wu et al. (2022) or fine-tuning on domain-specific corpora Liang et al. (2025). Ensuring that UrbanCIA agents reason within valid identification strategies, rather than relying on free-form generation, is essential for producing trustworthy and reproducible policy insights.
3.5 Evaluation
Evaluation Type Key Metrics Evaluators Representitives
Research Idea Novelty, Excitement, Feasibility, Effectiveness Human Si et al. (2024)
Significance, Clarity, Relevance, Originality, Feasibility Human, LLM ResearchAgent (2024)
Novelty, Value, Surprise, Relevance Human CoQuest (2024)
Method Development Clarity, Validity, Rigor, Innovativeness, Generalizability Human, LLM ResearchAgent (2024), MLR-Copilot (2024b)
Experiment Design Clarity, Validity, Robustness, Feasibility, Reproducibility Human, LLM ResearchAgent (2024)
Experiment Execution Performance Improvement, Success Rate Prototype Code MLR-Copilot (2024b)
Generated Paper Experimental Quality, Report Quality, Usefulness Human, LLM Agent Laboratory (2025)
NeurIPS-style Criteria Human, LLM The AI Scientist (2024)
Results Novelty, Impact Human, LLM AI co-scientist (2025)
Automated Reviewer Accuracy, F1 Score, AUC, FPR, FNR OpenReview Dataset (2022) The AI Scientist (2024)
Overall System Control, Creativity, Meta Creativity, Cognitive Load, Trust Human CoQuest (2024)
Utility, Continuation, Satisfaction, Usability Human Agent Laboratory (2025)
Cost (USD), Time (seconds), Success Rate System Logs Agent Laboratory (2025)
Elo Rating Human, LLM AI co-scientist (2025)
System’s Accuracy GPQA Dataset (2024) AI co-scientist (2025)
Safety Curated Adversarial Dataset AI co-scientist (2025)
Table 1: Overview of Evaluation Methods for Existing Autonomous LLM-Driven Research Works. Summary of evaluation types, key metrics, evaluators, and representative examples across research-related LLM-based works and autonomous research frameworks.
Evaluating the expert-level capabilities of LLM-driven systems is inherently challenging Si et al. (2024). The evaluation of UrbanCIA framework presents additional complexities, as it integrates multimodal data processing while maintaining the rigor required for causal inference experiments. Unlike general LLM-based automated research systems, which are primarily designed for machine learning tasks or scientific discovery in natural sciences, UrbanCIA introduces a novel paradigm tailored for urban causal research. This underscores the need to revisit the evaluation metrics for existing automated research systems, requiring evaluating their capacity to handle multimodal data, uphold the methodological rigor of causal inference, and ultimately generate meaningful causal insights that can inform urban policy and planning.
At present, evaluation in LLM-based research assistance or automated research encompasses a diverse range of quantitative and qualitative assessments, tailored to the specific tasks these systems perform. In Table 1, we present an organized overview of evaluation methods applied in autonomous LLM-driven systems, along with representative examples from recent studies. According to the overview of existing evaluation methods, LLM-driven systems are assessed not only on their ability to generate novel and feasible research ideas but also on their capacity to develop sound methodologies, design robust experiments, and execute research tasks effectively. These evaluations involve both human experts and automated systems or LLM-based reviewers, ensuring a balance between domain-specific judgment and scalable validation. For instance, research outputs (e.g., generated papers and reports), are scrutinized based on clarity, relevance, and scientific quality, often leveraging hybrid approaches that integrate peer review mechanisms with LLM-based assessments Si et al. (2024); Baek et al. (2024); Schmidgall et al. (2025); Lu et al. (2024).
Beyond evaluating individual research components, assessing the overall system performance is crucial to ensuring the reliability and ethical deployment of LLM-driven research systems. This evaluation encompasses multiple dimensions, including usability, interpretability, computational cost, and ethical considerations Liu et al. (2024); Schmidgall et al. (2025). In high-stakes decision-making contexts, particular emphasis is placed on safety to prevent potential misuse, such as dangerous research goals, dual-use risks, or misleading scientific claims Gottweis et al. (2025). One effective approach to system evaluation is testing performance in structured problem settings, such as benchmark datasetsBerto (2022); Lu et al. (2024) or curated adversarial datasetsGottweis et al. (2025).
For UrbanCIA, in addition to the existing evaluation methods, its inherent nature requires several additional evaluation focuses. We highlight three critical aspects: hypothesis generation significance, data processing quality, experiment design and causal inference validity. First, for idea generation evaluation, in addition to novelty and feasibility, it should also consider scientific rigor (i.e., ensuring asking appropriately clear and specific research questions adheres to established causal principles Bailey et al. (2024)) and interdisciplinary value (i.e., assessing whether they integrate insights from multiple domains, such as urban science, economics, and environmental studies, to provide innovative perspectives). Second, for data processing quality, its evaluation can focus on data completeness, consistency, preprocessing accuracy, scalability, and computational efficiency, ensuring robust, efficient, and effective multimodal data integration. Third, experiment design requires assessing the system’s ability to select appropriate causal identification strategies (e.g., DiD, IV), justify assumptions, and choose valid baselines. One practical evaluation approach is to reproduce classic urban causal inference studies and compare whether the agent’s design choices align with or diverge from established methodological best practices. Lastly, causal inference evaluation assesses causal structure feasibility, intervention validity, counterfactual consistency, and robustness to confounding. This can be achieved through synthetic data benchmarking, expert validation, or curated datasets derived from existing urban causal inference studies, ensuring that discovered causal insights are both statistically sound and practically relevant for urban policy applications. Table 2 summarizes the full set of evaluation dimensions and highlights the ones that are particularly specific to the UrbanCIA setting.
Evaluation Focus Key Aspects Specific?
Hypothesis Quality Causal clarity, scientific rigor, cross-domain relevance
Data Processing Completeness, temporal consistency, multimodal fusion quality
Causal Inference Rigor Intervention soundness, robustness checks, counterfactual consistency
Experiment Design Appropriate method choice, identification strategy, baseline comparisons
Reproducibility & Transparency Log traceability, reproducible pipeline setup, transparent assumptions
Ethical & Social Considerations Fairness, bias detection, value alignment, responsible intervention suggestion
System Usability & Efficiency Latency, scalability, agent coordination efficiency, task success rate
Table 2: Evaluation dimensions for UrbanCIA. The final column marks criteria that are particularly specific to UrbanCIA.
4 Further Discussion
In this section, we reflect on the broader implications of automating urban causal workflows, organized around three key perspectives: social impact, the human-AI collaboration paradigm, and ethical considerations. An overview of these dimensions is illustrated in Figure 6.
Figure 6: Overview of key perspectives on further discussion.. Three dimensions that shape the broader societal implications of AI-supported causal workflows in urban studies: social impact (who benefits), human-AI collaboration paradigms (how they interact), and ethical considerations (what concerns arise).
4.1 Social Impact
The proposed framework has the potential to transform scientific inquiry, policymaking, and public engagement by making causal insights more accessible, scalable, and actionable, as detailed below.
- Assisting Urban Researchers.
For urban researchers, UrbanCIA serves as a powerful assistant in accelerating and refining causal analysis. This allows researchers to focus on higher-level theoretical development and policy interpretations rather than labor-intensive preprocessing and computation. Moreover, UrbanCIA facilitates cross-city comparisons and large-scale causal experiments, enabling researchers to analyze urban dynamics across diverse spatial and temporal contexts. This can lead to more robust, generalizable findings that inform evidence-based urban policies and interdisciplinary research collaborations. Additionally, by standardizing causal inference workflows, UrbanCIA helps ensure reproducibility and transparency in urban studies, fostering a more rigorous and scalable approach to urban causal research.
- Enhancing Evidence-Based Urban Policy.
UrbanCIA provides policymakers with real-time causal analysis, allowing them to evaluate the potential impact of urban policies before implementation. By reducing reliance on correlation-driven decision-making, this automation ensures that interventions, such as traffic regulations, housing policies, or environmental protections, are based on causal insights rather than speculation. This can lead to more effective, data-driven, and proactive governance, ultimately improving urban livability, sustainability, and resource allocation.
- Democratizing Urban Analytics.
In addition, UrbanCIA lowers the barriers for citizens, journalists, and grassroots organizations to explore urban issues without requiring deep technical expertise. This fosters greater civic engagement and public awareness of how cities function. Urban studies should not only seek to explain the city but also be accessible to the people who live in it —because cities can only provide something for everybody when they are created by everybody Jacobs (1961). This ensures that data-driven insights contribute to more inclusive and participatory urban decision-making Fox and Wolf (2024).
4.2 New Human-AI Collaboration Paradigm
- From Tool to Collaborator.
Traditionally, AI systems have served as passive tools in urban research workflows, used for predictive modeling Li et al. (2024c), spatial visualization Zhang et al. (2024a), or data preprocessing Liang et al. (2025). However, the introduction of UrbanCIA represents a shift toward a more collaborative paradigm. In this framework, AI agents are no longer limited to executing predefined tasks; instead, they actively participate in the scientific reasoning process. Rather than merely accelerating isolated stages of analysis, these agents engage in context-aware interactions that reflect an evolving role: from tool to collaborator. This shift invites deeper questions about how tasks and responsibilities are distributed across human–agent systems, detailed as below.
- Division of Reasoning and Execution.
Moving beyond the traditional pipeline of human-led design, UrbanCIA embraces a configurable division of labor between humans and agents. Recent systems Gottweis et al. (2025); Li et al. (2024b); Lu et al. (2024); Swanson et al. (2024) reveal a common design principle: human experts define goals, constraints, and evaluation criteria, while agents handle iterative generation, optimization, and evidence organization. In these systems, natural language interfaces serve as the medium for task specification, correction, and refinement, reinforcing the human’s role as an agenda-setter and critical reviewer. In UrbanCIA, this flexible structure allows agents to operate independently for routine tasks such as data preprocessing or method selection, while preserving human judgment for hypothesis formulation, identification strategy, and assumption checking. Multi-agent mechanisms such as self-critique and ranking Lu et al. (2024) can further enhance robustness, but the final authority remains with human users, who validate outputs and determine their scientific and policy relevance. Rather than replacing human reasoning, this structure reinforces it: agents scaffold the causal workflow, while humans remain responsible for framing questions, defining standards, and interpreting outcomes in context.
- Language as Interface for Scientific Interaction.
Language interfaces enable more flexible and inclusive forms of human–AI collaboration in scientific reasoning. Recent systems Zhou et al. (2022); Swanson et al. (2024); Gottweis et al. (2025) demonstrate that researchers can use natural language to specify goals, refine hypotheses, and guide iterative experimentation without relying on formal task descriptions. This interaction model is especially valuable in urban causal research, where questions often involve loosely defined objectives and policy-relevant semantics. Through dialogic interfaces, users can engage in multi-round exchanges to clarify assumptions, compare methods, and request alternative designs. These interactions support asymmetrical collaboration, bridging technical experts, planners, and policymakers, and allow human judgment to remain embedded throughout the workflow. Moreover, by lowering technical barriers, such interfaces expand participation to civic actors such as journalists and community organizations, making urban inquiry more accessible and democratic Fox and Wolf (2024). Language thus serves not only as an input modality, but as a reasoning interface, shaping how tasks are interpreted, decisions are justified, and causal insights are co-constructed.
- Toward a New Paradigm.
Taken together, these shifts point toward a broader rethinking of how scientific inquiry is conducted. In UrbanCIA, causal analysis is no longer a static pipeline but a collaborative process, structured by agents, guided by human judgment, and mediated through language. As human and machine capabilities become increasingly entangled, urban research must move beyond automation toward shared reasoning systems that are adaptive, interpretable, and participatory.
4.3 Ethical Considerations
As AI agents take on active roles across the causal workflow, from hypothesis generation to policy-facing outputs, new ethical considerations arise around responsibility and transparency, fairness and bias, and the role of human oversight. First, when automated inferences inform public decisions, it is essential to clarify who is accountable for the assumptions, methods, and consequences. Human-in-the-loop designs can offer oversight, but they must be structured to avoid automation bias and ensure that AI-generated outputs remain contestable. To prevent misuse of AI tools by powerful corporations or institutions for their own advantage, interpretability should enable not only technical auditing but also meaningful engagement by domain experts and affected communities. Second, urban datasets often encode structural inequities. If agents are trained on dominant narratives or skewed data, they may inadvertently reinforce existing biases. Systems like UrbanCIA must embed mechanisms for uncertainty disclosure, sensitivity analysis, and inclusive data practices to mitigate this risk. Third, as AI becomes more embedded in scientific workflows, it raises broader questions around authorship, attribution, and epistemic authority. Rather than replacing human researchers, UrbanCIA is designed to support more inclusive, rigorous, and reflexive forms of urban causal reasoning—where domain expertise, civic participation, and algorithmic assistance can co-exist productively.
5 Summary & the Road Ahead
Reimagining Urban Causal Inference.
This Perspective highlights the untapped potential of LLMs to scale urban causal research. Urban causal research plays a vital role in evidence-based policymaking, yet current workflows are often labor-intensive, fragmented, and inaccessible. Our systematic review reveals five key gaps: geographic concentration, overreliance on structured data, limited methods, poor transparency, and slow causal adoption. To address these, we introduces UrbanCIA, a modular, LLM-powered framework that supports every stage of the causal pipeline, from hypothesis generation to evaluation, with an example use case. By structuring the process into discrete, automatable tasks, UrbanCIA show how AI can help scale causal research while retaining scientific rigor.
Lowering Barriers, Broadening Participation.
UrbanCIA is more than a tool. Instead, it is a step toward a more inclusive and participatory paradigm of urban inquiry. It lowers the technical threshold for causal analysis, enabling researchers, policymakers, and civic groups to explore timely urban questions with reduced dependence on coding, manual data handling, or advanced econometrics. In doing so, it fosters greater transparency, reproducibility, and responsiveness in how cities learn and adapt.
The Road Ahead.
Moving forward, we envision three key directions for future work. First, we plan to develop a deployable prototype of UrbanCIA to validate its effectiveness in real-world causal research workflows. Second, we aim to refine agent capabilities, especially for handling unstructured and multimodal data and designing rigorous causal inference experiments. Third, we will explore broader deployment scenarios in collaboration with urban researchers and policymakers. We call for collective efforts across AI, urban studies, and broader quantitative social science communities to realize a more inclusive, scalable, and rigorous future for urban causal inference.
Appendix A Related Works
A.1 Urban Causal Research
Urban causal research aims to identify the effects of treatments, policies, and interventions on outcome variables Baum-Snow and Ferreira (2015), which is crucial for evidence-based policymaking in areas such as transportation, housing, environmental sustainability, and economic development Baum-Snow and Ferreira (2015); Bailey et al. (2024). Causal inference methods (Hernan and Robins, 2020) thus have become indispensable tools in urban research, enabling more rigorous and data-driven policy evaluations or urban phenomenon analysis.
Randomized Controlled Trials (RCTs) are one of the commonly-used approaches for causal inference, effectively eliminating bias through random assignment. For example,  Bloom et al. (2024) employ RCTs to evaluate the impact of hybrid work-from-home policies on employee retention. However, RCTs are often impractical in urban research due to ethical, logistical, and spatial constraints. As a result, researchers rely more on quasi-experimental methods to infer causal relationships from observational data. One widely used approach is Difference-in-Differences (DID), which compares outcome changes between treated and control groups over time. This method has been applied to assess the impact of high-speed rail on traffic emissions, micromobility regulations on travel behavior, and vaccine mandates on public health Lin et al. (2021). Another key method is the Synthetic Control Method (SCM), which constructs a synthetic control unit as a weighted combination of untreated units that closely resemble the treated unit before intervention. SCM has been employed to evaluate the effects of COVID-19 lockdown policies on public sentiment and economic activity Wang et al. (2022). To address endogeneity issues, Instrumental Variable (IV) methods use external instruments that influence treatment but not the outcome directly. For example,  Zheng et al. (2024b) have leveraged the proportion of remote-capable workers to examine the impact of remote work on vehicle miles traveled and transit ridership. In addition, Propensity Score Matching (PSM) is used to reduce selection bias by creating comparable groups based on observable characteristics, and has been applied to evaluate the economic effects of electric vehicle charging stations Zheng et al. (2024a). Recently, with the advancement of Machine Learning (ML), ML-based causal inference methods, such as causal forests Ronco et al. (2023); Ito et al. (2024), have also gained popularity. In addition, deep neural networks for counterfactual inference Shalit et al. (2017); Louizos et al. (2017); Shi et al. (2019), along with their recent extensions to graph-structured data Wein et al. (2021), have further expanded the toolkit for learning complex causal effects in high-dimensional settings.
However, conducting robust urban causal research requires integrating domain expertise, high-quality data, and rigorous methodologies, making it a non-trivial task. From the data perspective, the high-volume, multi-source, and multi-modal nature of urban data demands extensive and repetitive human effort for data collection and pre-processing, reducing efficiency. In addition, causal methods rely heavily on subjective decisions in hypothesis formulation, variable selection, and methodological choices, which can introduce biases, overlook critical variables, and limit reproducibility.
A.2 Multimodal Large Language Models
Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) have recently emerged as powerful tools that process and generate human-like text by leveraging vast data and transformer-based architectures Vaswani (2017). Multimodal Large Language Models (MLLMs) extend the capabilities of LLMs by integrating and processing multiple data modalities, including text, images, audio, and structured data Wu et al. (2023). Unlike conventional LLMs, which primarily rely on textual inputs, MLLMs leverage cross-modal reasoning to enhance understanding and decision-making in complex tasks. In other words, MLLMs can be recognized as “universal translators” that bridge language and sensory inputs. Most MLLMs consist of three core components: a modality encoder that extracts features from non-text inputs, an LLM module that processes information, and a projector that aligns embeddings between different modalities Fu et al. (2024b). These models generate responses in an autoregressive manner, incorporating both textual and non-textual cues.
MLLMs have shown strong performance in tasks such as video captioning  Huang et al. (2024a); Fu et al. (2024a) and Visual Question Answering (VQA) Guo et al. (2023); Kuang et al. (2024). In the context of urban studies, MLLMs offer the potential to analyze multimodal urban data (satellite images, traffic sensors, social media), facilitating a more holistic understanding of urban dynamics Zhang et al. (2024c). For example, UrbanCLIP Yan et al. (2024) aligns satellite images with textual descriptors to profile a certain urban region. In transportation, real-time traffic data combined with social media mentions of a concert allows systems such as TrafficGPT Zhang et al. (2024a) to diagnose congestion triggers and optimize routing. For disaster resilience, DisasterResponseGPT Goecks and Waytowich (2023) dynamically generates emergency plans by contextualizing real-time data against crisis protocols.
A.3 AI for Scientific Discovery
Recent advancements in AI have driven a paradigm shift in scientific discovery. Specifically, AI has demonstrated its potential in scientific breakthroughs, such as AlphaFold 2 Jumper et al. (2021) in protein structure prediction, advancing drug discovery and materials science. Beyond specialized models, recent efforts integrate LLMs into the entire research workflow, from hypothesis generation to manuscript writing. Studies suggest LLM-generated feedback can be as effective as human peer reviews Liang et al. (2024). Several AI-driven systems aim to assist in different stages of scientific research: PaperQA Skarlinski et al. (2024) aids literature search, while HypoGeniC Zhou et al. (2022) facilitates hypothesis generation through iterative refinement and platforms like data-to-paper Ifargan et al. (2025) and Virtual Lab Swanson et al. (2024) employ LLM-based agents to automate research paper writing and experimental design. Multi-agent systems like Coscientist Boiko et al. (2023) and The AI Scientist Lu et al. (2024) extend this paradigm by autonomously handling hypothesis formulation, experimental execution, and research documentation. Recently, the AI co-scientist Gottweis et al. (2025), a multi-agent system built on Gemini 2.0, demonstrated its potential in drug repurposing, novel target discovery, and bacterial evolution.
Though promising, most AI-driven scientific discovery primarily focused on natural science Yan et al. (2025); its application to social science research remains in its early stages. Unlike the physical and life sciences, where AI can generate and validate hypotheses through controlled experiments, social science research involves complex, evolving human behaviors that are harder to formalize and test empirically. Recent efforts have begun addressing this gap. AI-assisted frameworks have been developed to uncover quantitative and symbolic models, bridging parametric and non-parametric approaches Balla et al. (2025).  Manning et al. (2024) leverages structural causal models, where LLMs autonomously generate and test social science hypotheses through in silico experiments. Additionally, LLMs have been applied to social skill training and human-AI interaction, demonstrating potential in communication research and behavioral science Yang et al. (2024a). More recently, LLM-driven simulations model human behavior and societal dynamics, enabling controlled studies on polarization, misinformation, and economic policies Piao et al. (2025b, a). However, AI-driven discovery in urban causal research remains largely unexplore.
Appendix B Formulations
B.1 Causal Inference Problem
In a fundamental causal inference problem, the objective is to estimate the causal relationship between a set of treatment variables $T$ and an outcome $y$. For each observation $i$, the data generation process of outcome $y_{i}$ can be denoted as follows Baum-Snow and Ferreira (2015)
$y_{i}=T_{i}\beta_{i}+X_{i}\delta_{i}+U_{i}+e_{i},$ (1)
where $y_{i}$ represents the observed outcome for unit $i$, influenced by a treatment variable $T_{i}$, a set of observed control variables $X_{i}$, unobserved confounders $U_{i}$, and a stochastic error term $e_{i}$. The coefficients $\beta_{i}$ and $\delta_{i}$ capture the respective effects of the treatment and control variables on the outcome. In the context of urban studies, the outcome variable $y_{i}$ can represent an urban phenomenon, e.g., air pollution levels, traffic congestion, housing prices at a given location $i$. The treatment variable $T_{i}$ typically corresponds to an urban policy intervention or infrastructure change, such as the implementation of congestion pricing, new public transit infrastructure, or zoning regulation modifications.
The primary objective is to estimate $\beta_{i}$, which quantifies the causal effect of the treatment on the outcome while controlling for confounders $X_{i}$. Various econometric and statistical methods can be employed to estimate these coefficients, including but not limited to Ordinary Least Squares (OLS) for simple regression-based analysis, Difference-in-Differences (DiD) for policy impact evaluation, Propensity Score Matching (PSM) for balancing covariates, Instrumental Variables (IV) to address endogeneity, and Synthetic Control Methods (SCM) for counterfactual estimation.
B.2 LLMs & MLLMs
Large Language Models (LLMs) Achiam et al. (2023); Touvron et al. (2023); Team et al. (2023); Guo et al. (2025) are trained on large-scale text corpora to predict token sequences in an autoregressive manner. Given a sequence of tokens $\mathbf{x}=(x_{1},x_{2},\dots,x_{t})$, an LLM models the conditional probability of the next token as:
$p(x_{t}\mid\mathbf{x}_{<t};\theta)=\text{softmax}(\mathbf{W}\cdot\mathbf{h}_{t}),$ (2)
where $\mathbf{h}_{t}$ denotes the hidden state at time step $t$, and $\mathbf{W}$ is the learned weight matrix mapping hidden states to token logits. The model generates text by sampling iteratively from this distribution.
Multimodal Large Language Models (MLLMs) extend this paradigm by incorporating additional modalities, e.g., imagery Liu et al. (2023a); Yan et al. (2024), audio Chu et al. (2023), graph Tang et al. (2024), and structured charts Han et al. (2023), alongside text. These non-text inputs are embedded into a shared representation space as $\mathbf{e}_{V}$, while the textual prompt is embedded as $\mathbf{e}_{T}$. The MLLM then generates output tokens $\mathbf{y}=(y_{1},y_{2},\dots,y_{L})$ according to:
$p(\mathbf{y}\mid\mathbf{e}_{V},\mathbf{e}_{T})=\prod_{t=1}^{L}P(y_{t}\mid\mathbf{y}_{<t},\mathbf{e}_{V},\mathbf{e}_{T}),$ (3)
where $\mathbf{y}_{<t}$ denotes previously generated tokens.
By integrating multimodal contexts, MLLMs significantly expand the capabilities of LLMs—enabling tasks such as image captioning, visual question answering, and multimodal reasoning that are beyond the scope of text-only models.
Symbol Description
$\mathbf{x}=(x_{1},x_{2},\dots,x_{t})$ Input sequence of textual tokens
$x_{t}$ Token at time step $t$
$\mathbf{x}_{<t}$ Sequence of tokens preceding time step $t$
$\mathbf{h}_{t}$ Hidden state of the LLM at step $t$
$\mathbf{W}$ Weight matrix mapping hidden states to token logits
$\mathbf{e}_{V}$ Embeddings from non-text modalities (e.g., images, audio)
$\mathbf{e}_{T}$ Textual input embeddings
$\mathbf{y}=(y_{1},y_{2},\dots,y_{L})$ Generated output token sequence of length $L$
$p(x_{t}\mid\mathbf{x}_{<t};\theta)$ Conditional probability of token $x_{t}$ given history
$P(y_{t}\mid\mathbf{y}_{<t},\mathbf{e}_{V},\mathbf{e}_{T})$ Conditional probability of multimodal token generation
Table 3: Notation. Notation used for LLMs and MLLMs formulation.
Appendix C Methodology for Systematic Review
In Section 2, we present a comprehensive systematic review that maps the current landscape of causal inference in urban studies. The objective is to obtain a quantitative understanding of the distribution of urban causal inference studies across five key dimensions: research topic, geographic focus, data modality, experiment type, and methodological approach. Here we detail the methodology used to conduct the review as follows. The prompts used for each step and the code implementing this procedure can be found at: https://github.com/quao627/Urban-CI-LitReview.
Step 1: Identify Papers That Conduct Causal Inference Studies
Among the 2,428 papers downloaded from the Cities journal, we identify those conducting causal inference studies using a crafted prompt with the GPT-4o model. This filtering process involves two steps. First, we exclude papers that are clearly out of scope by examining only their titles and abstracts. Second, the LLM accesses the full text of the remaining papers and labels each one accordingly. This two-step process helps reduce costs by eliminating irrelevant papers early using a low-cost method.
Step 2: Paper Parsing via LLMs.
We develop a structured prompt to guide GPT through the analysis of each paper, with the aim of extracting relevant information and supporting evidence corresponding to each dimension. To retrieve content from full-text PDFs, we utilize the OpenAI API’s file upload feature [^5] https://platform.openai.com/docs/guides/pdf-files?api-mode=chat, which allows direct querying over document contents. This step produced initial annotations per dimension for every paper.
Step 3: Iterative Label Refinement.
For each dimension, we design an iterative clustering pipeline that leverages GPT-4o to semantically group similar entries and propose representative labels. The algorithm identifies recurring patterns, merges equivalent concepts, and refines the terminology. This step outputs a set of standardized keywords describing each paper across the five dimensions.
Step 4: Human Validation.
To ensure accuracy and domain relevance, we manually review and refine the final set of clusters and labels using expert judgment. This final validated dataset serves as the basis for the visualizations and quantitative analyses presented in the main text.
Appendix D Roles and Functions of Agents in the Example Use Case
This section outlines the roles and responsibilities of each agent involved in the NYC congestion pricing use case illustrated in Figure 5.
  • Reader. Continuously gathers multimodal information from external sources (e.g., policy announcements, social media, academic papers) to identify emerging issues and formulate preliminary hypotheses.
  • Urban Scientist. Refines hypotheses by incorporating domain expertise (e.g., local infrastructure, demographics, governance context) and evaluates their feasibility and policy relevance to ensure that the research question is well-scoped and actionable. It also suggests relevant data variables required for investigation.
  • Data Engineer. Builds the data pipeline to collect, process, and align multimodal urban data (e.g., visual, textual, temporal, spatial). Responsibilities include data cleaning, missing value imputation, and spatio-temporal normalization across sources.
  • Validator. Oversees quality control at multiple stages: (1) verifies data readiness (e.g., treatment/control group coverage, covariate completeness), (2) assesses robustness of results (e.g., placebo tests, sensitivity checks), and (3) ensures narrative coherence and policy alignment of the final report.
  • Experimenter. Implements the experimental design by executing data analysis tasks (e.g., exploratory regressions, causal estimators) and generating outputs such as subgroup estimates and summary visualizations.
  • Data Scientist. Designs the causal inference strategy by selecting appropriate identification methods (e.g., quasi-experiments, structural models), defining confounders and assumptions, and guiding the implementation plan.
  • Writer. Integrates analytical findings and policy interpretations into a coherent report. It composes textual explanations, curates visual summaries, and ensures the final output communicates insights that are both rigorous and decision-relevant.